Leia is considering insuring against theft the new $300 CD-player she just installed in her automobile.
Her insurance agent tells her that such an option could be added to her present policy for $20 per year.
The agent further states that the probability of theft is 0.2 in a given year. If she takes the insurance,
what is her expected return per year?
1
Expert's answer
2020-11-18T18:58:41-0500
"Solution"

Cost of CD-player = "\$300"

Cost of additional insurance = "\$20" per year

Probability of theft is = "0.2"


"\therefore" probability of no theft = "1 - 0.2 = 0.8"

If there is no theft, then amount paid to insurance is "-\$20"

If there is theft, then amount paid to insurance is "\$300-\$20=\$280" which is (cost of CD-player recovered - Cost of additional insurance)


So expected return per year

"\sum P_i \times r_i" where "P_i" is the probability of event(theft or no theft), and is the "r_i" return amount


"0.8 \times (-20\$)\ +\ 0.2 \times \$280 = \$40"

Hence the expected return per year is "\$40"